please help me solve this word problem. A model house has a scale of 2 inches : 3 feet. if the real house is 28 feet wide, then how wide is the model house?
Answer by stanbon(75887)

A model house has a scale of 2 inches : 3 feet. if the real house is 28 feet wide, then how wide is the model house?
----------------
model/real = model/real
-----
model/28 = 2 in/3 ft.
--------
model = 28(2/3) in = 56/3 = 18 3/4 inches
====================
Cheers,
Stan H.
====================